The normal to the parabola y2= 4ax at the point P(at2 , 2at) meets the x-axis at A. Find the equation of the locus of the midpoint of AP as t varies.


1
Expert's answer
2021-08-26T14:31:53-0400

Solution:

To find the normal;

At A(y,x) but y=0

Since;

"y^2=4ax"

"\Rightarrow 2y\frac{dy}{dx}=4a"

"\Rightarrow \frac{dy}{dx}=\frac{2a}y"

Slope of the tangent at

"P(at^2,2at)" ;"m_1=\frac{dy}{dx}=\frac{2a}{2at}" ="\frac1t"

Hence slope of the normal at P is given as;

"m_2=\dfrac{-1}{\dfrac1t}" ="-t"

Therefore equation of the normal is given by;

"y=m_2x+c
\\\Rightarrow y=-tx+c"

Put "x=at^2,y=2at"

So, "c=2at+at^3"

Thus, "tx+y=2at+at^3"

Hence;

"y-2at=-t(x-at^2)"

But y=0

"-2at=-tx+at^3"

Rewrite as;

"tx=at^3+2at"

Divide by t,we get;

"x=at^2+2a"

Take the midpoint of AP to be B(h,k)

Here;

"h=\frac{at^2+2a+at^2}{2}" ="at^2+a"

"k=\frac{0+2at}{2}" ="at"

Since ;

"y^2=4ax"

We have the equation of the locus as t varies given by at P;

"(2at)^2=4a(at^2)"

Using h and k ,we can make the following substitutions;

"(2\u00d7k)^2=4a(h-a)"

"\Rightarrow 4k^2=4ah-4a^2"

Rewrite as ;

"k^2=ah-a^2"

"\Rightarrow k^2-ah+a^2=0"

Writing it in terms of x and y as;

"y^2-ax+a^2=0" which is the standard locus at midpoint AP.